OpenStudy (anonymous):

The drama club at Lincoln High School sells hot chocolate and coffee at the school's football games to make money for a special trip. At one game, they sold $194.50 worth of hot drinks. They need to report how many of each type of drink they sold for their club records. Macha knows that they used 286 cups that night. If hot chocolate sells for 75 and coffee sells for 50, how many of each type of hot drinks did they sell?

OpenStudy (anonymous):

Picture please

OpenStudy (anonymous):

i dont have a picture thats the question

OpenStudy (anonymous):

x+y=286 ---> x=286-y 0.75x+0.5y=194.5 --------------- 0.75(286-y)+0.5y=194.5 214.5-0.75y+0.5y=194.5 -0.75y+0.5y=194.5-214.5 -0.25y=-20 y=20/0.25 y=80 --------------- x+y=286 x=286-y x=286-80 x=206 they sold 206 cups of chocolate & 80 cups coffee let's check 206*0.75=154.5 80*0.5=40 40+154.5=194.5

OpenStudy (anonymous):

Another approach using Mathematica 9 for the calculations. Refer to the attachment.